# Feasibility Study of CNNs and MLPs for Radiation Heat Transfer in 2-D Furnaces with Spectrally Participative Gases

### v2

@@ 3.2 CNN
As can be seen, each input parameter (emissivity, boundary temperature, domain temperature) is considered as a channel. We just set the values to zero for those that don’t exists (comments: no s), e.g. wall temperature and emissivity in the domain, domain temperature on the wall. The output layer is the same as the one used with MLP architecture.
@@ 4 Results and discussions
By comparing these values, one can see that the CNN yields better results than the MLP, still the relative error for both cases are high (comment: errors or is). One can also observe the effect of larger datasets on the training of both networks. For MLP, the larger dataset B improved the mean and STD values by $10.14\%$ and $0.34\%$, respectively. Similarly, for CNN, the mean and STD values are ameliorated by $18.41\%$ and $15.76\%$, respectively.
The relative errors for each network are shown in Fig. 4. The points on the horizontal line (comment: with s or is):
@@ 5 Summary
The results showed that both MLP and CNN can indeed learn the solution of RTE. Compared to MLP, CNN performed better in terms of accuracy and speedup. For the cases when the output of both networks corresponded to all the boundary points, the mean and standard deviation of the relative error of CNN was $9.07\%$ and $1.84\%$, respectively. Also, the speedup was about x11000. We also saw that larger dataset improved the accuracy of both networks. However, the improvement was bigger for CNN network than MLP network ($18\%$ v.s. $10\%$). Then, we chose the output vector to correspond to only one boundary wall. Having kept the same hyper-parameters, we observed that the new reduced-in-size output vector for only one boundary wall enhanced the accuracy of CNNs but not the MLPs.(comment: This) this means that the CNN outperformed the MLP in terms of robustness and stability against hyper-parameter changes.

### v3

@@ 3.2 CNN
As can be seen, each input parameter (emissivity, boundary temperature, domain temperature) is considered as a channel. We just set the values to zero for those that don’t exist, e.g. wall temperature and emissivity in the domain, domain temperature on the wall. The output layer is the same as the one used with MLP architecture.
@@ 4 Results and discussions
By comparing these values, one can see that the CNN yields better results than the MLP, still the relative errors for both cases are high. One can also observe the effect of larger datasets on the training of both networks. For MLP, the larger dataset B improved the mean and STD values by $10.14\%$ and $0.34\%$, respectively. Similarly, for CNN, the mean and STD values are ameliorated by $18.41\%$ and $15.76\%$, respectively.
The relative errors for each network are shown in Fig. 4. The points on the horizontal axis:
@@ 5 Summary
The results showed that both MLP and CNN can indeed learn the solution of RTE. Compared to MLP, CNN performed better in terms of accuracy and speedup. For the cases when the output of both networks corresponded to all the boundary points, the mean and standard deviation of the relative error of CNN was $9.07\%$ and $1.84\%$, respectively. Also, the speedup was about x11000. We also saw that larger dataset improved the accuracy of both networks. However, the improvement was bigger for CNN network than MLP network ($18\%$ v.s. $10\%$). Then, we chose the output vector to correspond to only one boundary wall. Having kept the same hyper-parameters, we observed that the new reduced-in-size output vector for only one boundary wall enhanced the accuracy of CNNs but not the MLPs. This means that the CNN outperformed the MLP in terms of robustness and stability against hyper-parameter changes.